# Non-Relativistic Hyperboloidal Method for Quasinormal Modes

### v1

@@ I Introduction
Numerical methods underpin much of the progress in QNMs over recent years. Indeed, efficient schemes for computing the QNMs of potentials are likely indispensable for future developments in both theory and the modelling of observations. Recently, the so-called compactified hyperboloidal method has proven to be a powerful tool Jaramillo2021 , finding wide use in the computation of black hole QNM spectra and bringing within reach the systematic exploration of their connection to pseudospectra Destounis2021 ; Ripley2022 ; Gasperin2022 ; Sarkar2023 ; Boyanov2024 ; Cao2024 ; Zhu2024 ; Destounis2024 . Beyond this, it is natural to ask whether the method can also find use in optical systems. We believe it can, but it cannot be widely applied in optics without modification. This is because optical media create non-relativistic and dispersive dynamics, while the present formulation of the method treats only relativistic and non-dispersive dynamics, as may be seen from its use of hyperbolic spatial slices penetrating the black hole horizon and future null infinity.
A notable optical system that motivates the development of a hyperboloidal method for optics is the fiber optical soliton, which has recently been established as a black hole analogue with an exactly known QNM spectrum Burgess2024 . As such, the soliton is the ideal system with which to develop the method, as the resulting numerics can be compared both to known analytical results and to the numerics of the corresponding relativistic system. Moreover, perturbations to the soliton realize the Schrödinger equation with a Pöschl-Teller potential, making the soliton a promising experimental platform with which to address questions in QNM theory, such as the physical status of spectral instabilities observed in QNM numerics, where the Pöschl-Teller potential is paradigmatic Ferrari1984 ; Nollert1993 ; Cho2012 ; Jaramillo2021 ; Cardoso2024 .
@@ II Compactified Hyperboloidal Method for the Schrödinger Equation
with $V$ a potential that vanishes for $r\to\pm\infty$. In order to compute the QNMs of this equation, we first transform from coordinates for which the QNMs diverge at the spatial boundaries, to ones for which they remain finite. Following Jaramillo2021 , we adopt coordinates,
@@
where $g(x),h(x)$ are yet to be given, and $\partial_{\tau}=\partial_{t}$ by construction. In the relativistic method, $h(x)$ is chosen so that contours of $\tau$ tend asymptotically to the horizon and light cone. However, there is no preferred speed in our system, and we parameterize $h(x)$ to obtain hyperbolic slices propagating outwards with asymptotic speed $v_{a}$ in the $(r,t)$ coordinates. In particular, we write
$g(x)=\arctan{x},\quad h(x)=v_{a}^{-1}\log(1-x^{2}),$ (3)
where $g(x)$ compactifies the space such that the real line of $r$ gives $x\in[-1,1]$ if we close the set by including the boundaries. By considering mode solutions far from the potential, it is easily confirmed that those QNMs whose contours of constant amplitude propagate outwards with speed $v_{a}$ take finite amplitudes on the new spatial boundaries. However, dispersion in this non-relativistic system implies that asymptotic contours of constant amplitude do not generally align with those of constant phase, meaning finite amplitudes on the boundaries are typically accompanied by phase singularities. This is resolved by a phase-rotation of the field,
@@
with $\Delta=i(v_{a}-v_{p})/2$ and $v_{p}$ the asymptotic phase velocity associated with $\phi$. The form of the phase rotation follows from the asymptotic dispersion relation of Eq. (1), and intuitively depends on the mismatch of the two speeds. This ensures $\hat{\phi}$ is well-defined on the compactified space, but its construction presumes two real parameters, $v_{a}$ and $v_{p}$. Indeed, for mode solutions, these parameters are characteristic, in that they determine the frequency of the mode. In particular, we find
$\Omega=-\frac{1}{2}v_{a}\bigg{(}v_{p}+i\sqrt{v_{a}(v_{a}-2v_{p})}\bigg{)},$ (5)
having used $\Omega=-K^{2}$, the asymptotic dispersion relation of Eq. (1), where $K$ is the asymptotic wavenumber in the $(t,r)$ coordinates. That we can express the mode frequency in terms of asymptotic speeds is a direct consequence of dispersion. In a relativistic system, all asymptotic speeds are the speed of light, and $\Omega$ cannot be expressed in terms of asymptotic speeds. This difference between the relativistic and non-relativistic methods is crucial, as will be explained in what follows.
@@
where $J_{1}$ and $J_{2}$, given in the Appendix, are spatial operators depending on the potential and the asymptotic speeds. In contrast to the relativistic method, $\partial_{\tau}\hat{\psi}$ cannot be isolated by division in Eq. (6) because its pre-factor vanishes at $x=0$. This unavoidably occurs as a consequence of the hyperboloidal slices necessarily “turning around” in order to describe contours that are outgoing in both directions. Our alternative approach is to construct the $\partial_{\tau}\hat{\psi}$ using a Taylor series of its spatial derivatives at $x=0$, which may be obtained by repeated differentiation of Eq. (6). In fact, this treatment is necessary only for those terms indivisible by $x^{2}$, and we obtain a simpler result if we initially separate the terms in this way. This separation is mostly trivial, but for the potential we write $V=V_{0}+xV_{1}+x^{2}\tilde{V}(x)$, where $V_{0}$ and $V_{1}$ are Taylor series coefficients about $x=0$ and $\tilde{V}$ accounts for the remaining terms. We obtain
@@
where $L_{1}$ and $L_{2}$ are spatial operators that we derive in the Appendix. Eq.(7) is formally identical to that obtained in the relativistic method Jaramillo2021 , but the operators are quite different, containing arbitrarily high spatial derivatives and depending on the asymptotic speeds.
@@
Importantly, the operator $L$ in the above equation is actually parameterized by the mode frequency $\Omega$. This is because the operator depends on $v_{a}$ and $v_{g}$, which may be expressed in terms of $\Omega$ and its complex conjugate using Eq.(5). This is an essential difference from the relativistic method, where the corresponding operator contains no dependence on $\Omega$ Jaramillo2021 . Nevertheless, Eq.(8) unambiguously determines the QNM spectrum.
@@
The QNM spectrum may then be obtained in the usual way using $\det(L^{N}-\Omega\ \!\text{Id})=0$. In the Appendix, we show that this determinant may be rewritten as that of a smaller $N$-dimensional matrix, $M$. Its elements are quadratic in the square root of the QNM frequency, giving rise to a polynomial of degree $2N$ in $\sqrt{\Omega}$. For a given potential $V$, the roots may be numerically determined in order to give $2N$ of the QNM frequencies. The fact that the frequency enters via its square root is a result of the Schrödinger equation having a first derivative in time, rather than a second derivative in time like the wave equation. Indeed, the exact QNM spectra of the Schrödinger and wave equations are related to each other by $i\sqrt{\Omega}=\omega$, as elaborated in Burgess2024 . This means we can relate the results of the non-relativistic method to those of the relativistic method, allowing us to better evaluate the accuracy of the new method.
@@ III Quasinormal Modes of the Pöschl-Teller Potential
which serves as an exemplar for both the relativistic and non-relativistic methods. Its QNMs are finite polynomials in the compactified spatial coordinate, with the result that an $N$-point discretization reproduces the first $2N$ QNMs to arbitrary precision. The Pöschl-Teller potential is also ideal because the corresponding QNM spectrum of the Schrödinger equation is given analytically by
@@
allowing us to verify our results Burgess2024 . In regards to the non-relativistic method, we note that the Pöschl-Teller potential is especially simple because all its QNMs have the same $v_{a}$ parameter, which is a result of the unique fact that $i\sqrt{\Omega}$ is aligned along vertical lines in the complex plane for this potential. Lastly, we partition the Pöschl-Teller potential with $V_{1}=0$ and $\tilde{V}=-V_{0}$, which reinforces the simplicity of the potential.
@@
Figure 1: QNM spectra for the Schrödinger equation with a potential $V=V_{0}\hskip 1.5pt\text{sech}^{2}(r)+\epsilon\Delta V$, where $V_{0}=1$ and $\Delta V$ is the perturbation shown in green in the inset, alongside the unperturbed Pöschl-Teller potential in blue. The red crosses and blue circles correspond to the unperturbed Pöschl-Teller potential ($\epsilon=0$), with red crosses given by the exact formula and blue circles numerically determined by the new method. The green diamonds correspond to a perturbed potential ($\epsilon=10^{-30}$) and are also numerically determined. (a) displays the three QNM spectra, while (b) displays the same spectra under the transformation $\Omega\to i\sqrt{\Omega}$, which relates the spectra to those of a corresponding relativistic operator.
In Figure 1(a), we plot the exact QNM frequencies of the unperturbed Pöschl-Teller potential Burgess2024 alongside those calculated by the new numerical method, with a resolution of $N=201$. We find excellent agreement for all frequencies, with an error which may be made arbitrarily small by increasing the working precision. These results are given in Figure 2. We also calculate the QNM spectrum of a perturbed Pöschl-Teller potential,
@@
where $\epsilon=10^{-30}$ and $\Delta V$ is a randomly chosen polynomial of degree 9, shown in the inset of Figure 1. We find that the spectrum closely resembles the unperturbed spectrum up to the $10^{\text{th}}$ overtone index, beyond which the frequencies are significantly displaced from their unperturbed values, as shown in Figure 2. These numerical results are then indicative of spectral instabilities that have been reported by previous authors Nollert1996 ; Daghigh2020 ; Jaramillo2021 .
Figure 2: Comparisons of exact and numerically determined QNM frequencies for the Schrödinger equation with a potential $V=V_{0}\hskip 1.5pt\text{sech}^{2}(r)+\epsilon\Delta V$, where $V_{0}=1$ and $\Delta V$ is the perturbation shown in Figure 1. The first 20 QNM frequencies are displayed. The unperturbed ($\epsilon=0$) spectrum is recovered well by the new numerical method, with errors smaller than $10^{-18}$ for the chosen working precision. We also obtain the perturbed ($\epsilon=10^{-30}$) spectrum and find the deviation from the exact spectrum grows radidly with overtone index, $n$, as in previous works on spectral instability.
@@ IV Discussion
In this section, we discuss potential applications of the non-relativistic compactified hyperboloidal method that we developed in the preceding text, suggesting well-motivated directions in which to further develop the method and providing a sketch of how this can be achieved. The principal application of the method is the calculation of QNM frequencies of the Schrödinger equation equipped with a general potential. Using our method, we numerically calculated QNM spectra for the Pöschl-Teller potential and perturbations of that potential, finding agreement with earlier works Boonserm2011 ; Cardona2017 ; Burgess2024 . In addition, we note that this method may also be used to numerically solve for the quantum mechanical bound-states of a general potential well, using the well-known connection between the QNMs of a potential barrier and the bound-states of the corresponding well Ferrari1984b .
@@
A further motivation for generalizing the non-relativistic method is to shed light on QNM spectral instabilities, and facilitate experimental tests of the recent ultraviolet universality conjecture, which posits that sufficiently high overtones converge to logarithmic Regge branches in the complex plane, in the high-frequency limit of potential perturbations Jaramillo2021 ; Boyanov2024 . This effect is easily seen in numerical calculations of the Pöschl-Teller spectrum, on account of its simplicity, but has yet to be experimentally confirmed. Using the optical soliton, whose perturbations realize this potential, an experimental test becomes possible. However, interpreting observations of the migration of soliton QNM frequencies demands an understanding of the higher-order dispersive effects which will be present in any experiment. In particular, the influence of weak third-order dispersion acting on the perturbative probe field should be incorporated into the analysis, in order to provide the best test of the above conjecture. This motivates the development of the non-relativistic method beyond the Schrödinger equation, to include higher-order dispersive terms.
@@
with $\alpha(z)$ and $\beta(z)$ finite polynomials in $z$, and $d$ the larger degree among the two polynomials. We apply a hyperboloidal coordinate transformation and a phase rotation of the fields, parameterised by the asymptotic speeds $v_{a}$ and $v_{p}$. Then, we introduce auxiliary fields to effect a $d$th-order reduction in time, defining
@@
with $1\leq k<d$. This closely mirrors the treatment of resonator QNMs in optics lalanne2019 . The equations of motion of these fields are trivial for all fields but $\phi_{d}$, whose equation of motion more closely resembles Eq. (6). If we use a Taylor series expansion of $\partial_{\tau}\phi_{d}$ around zero, we can write it in terms of spatial operators acting on the fields. The general form of the now $d$-dimensional operator $L$ is
@@
which we discretize as before. Then, we use the asymptotic dispersion relation of Eq. (13) to eliminate the asymptotic speeds, obtaining a vector equation for the QNM frequencies. From the form of $L$, it can be shown that it is always possible to construct an $N$-dimensional matrix $M$ whose determinant is a finite polynomial for the QNM frequencies. This may then be solved numerically and the frequencies $\Omega$ determined.
Future works can also develop this non-relativistic method, along the lines of Jaramillo2021 , in order to calculate the pseudospectra of non-relativistic operators. It is our view that the relationship between perturbed QNM spectra and the pseudospectrum is best understood from a broader perspective, not limited to relativistic wave operators. Lastly, we expect that numerical methods will become increasingly important for addressing questions in QNM theory, and we anticipate that investigations into the QNMs of non-relativistic fields will provide new avenues to explore these questions.
@@
We would like to express our thanks to Théo Torres for providing us a useful overview at the outset of this research. This work was supported in part by the Science and Technology Facilities Council through the UKRI Quantum Technologies for Fundamental Physics Programme [Grant ST/T005866/1]. CB was supported by the UK Engineering and Physical Sciences Research Council [Grant EP/T518062/1].
For the purpose of open access, the authors have applied a creative commons attribution (CC BY) licence (where permitted by UKRI, ‘open government licence’ or ‘creative commons attribution no-derivatives (CC BY-ND) licence’ may be stated instead) to any author accepted manuscript version arising.

### v4

@@ I Introduction
Numerical methods underpin much of the progress in QNMs over recent years. Indeed, efficient schemes for computing the QNMs of potentials are likely indispensable for future developments in both theory and the modelling of observations. Recently, the so-called compactified hyperboloidal method Zenginoglu2008 ; Zenginoglu2011 ; Jaramillo2021 ; Macedo2024 has proven to be a powerful tool, finding wide use in the computation of black hole QNM spectra and bringing within reach the systematic exploration of their connection to pseudospectra Destounis2021 ; Ripley2022 ; Gasperin2022 ; Sarkar2023 ; Boyanov2024 ; Cao2024 ; Zhu2024 ; Destounis2024 . Beyond this, it is natural to ask whether the method can also find use in optical systems. We believe it can, but it cannot be widely applied in optics without modification. This is because optical media create non-relativistic and dispersive dynamics, while the present formulation of the method treats only relativistic and non-dispersive dynamics, as may be seen from its use of hyperbolic spatial slices penetrating the black hole horizon and future null infinity.
A notable optical system that motivates the development of a hyperboloidal method for optics is the fiber optical soliton, which has recently been established as a black hole analogue with an exactly known QNM spectrum Burgess2024 . As such, the soliton is the ideal system with which to develop the method, as the resulting numerics can be compared both to known analytical results and to the numerics of the corresponding relativistic system. Moreover, perturbations to the soliton realize the Schrödinger equation with a Pöschl-Teller potential, making the soliton a promising experimental platform with which to address questions in QNM theory, such as the physical status of spectral instabilities observed in QNM numerics, where the Pöschl-Teller potential is paradigmatic Ferrari1984a ; Nollert1993 ; Cho2012 ; Jaramillo2021 ; Cardoso2024 .
@@ II Compactified Hyperboloidal Method for the Schrödinger Equation
with $V$ a potential that vanishes for $r\to\pm\infty$. The boundary conditions for QNMs describe solutions that transport energy away from the potential, as discussed in more detail in Burgess2024 . It can be shown, using the asymptotic dispersion relation of Eq. (1), that QNM solutions must diverge for $r\to\pm\infty$. That is, the asymptotic form of the solution must be $\phi\sim\exp(iKr-i\Omega t)$, with the requirement that $\text{Im}(K)$ is positive and negative on the left and right, respectively. These spatial divergences are problematic for numerical methods, but they can be removed by using a hyperboloidal coordinate transformation. Following Jaramillo2021 , we adopt coordinates,
@@
where $g(x),h(x)$ are yet to be given, and $\partial_{\tau}=\partial_{t}$ by construction. In the relativistic context, these are used to compute QNMs of black holes, with $h(x)$ chosen so that contours of $\tau$ tend to null curves that intersect the horizon and future null infinity. There, Eq. (2) is intended to respect the asymptotic hyperbolic geometry of the spacetime, giving rise to bounded and well-behaved QNM solutions. However, there is no preferred speed in our non-relativistic system, meaning that no coordinate transformation will consistently give rise to bounded solutions. This requires a different approach.
In order to construct bounded QNM solutions, we first parameterize $h(x)$ by a new variable $v_{g}$ such that contours of $\tau$ tend asymptotically to trajectories directed outwards with $|dr/dt|=v_{g}$. In particular, we write
$g(x)=\text{arctanh}\ \!{x},\quad h(x)=(2v_{g})^{-1}\log(1-x^{2}),$ (3)
where $g(x)$ compactifies the space such that the real line of $r$ gives $x\in[-1,1]$ if we close the set by including the boundaries. In the Appendix, we show that QNMs whose asymptotic group velocity is $v_{g}$ in $(r,t)$ coordinates are finite at the spatial boundaries, $x=\pm 1$. This enforces the boundary conditions for these modes, but does not guarantee that any such modes exist.
In contrast to the relativistic case, dispersion in non-relativistic systems means that group and phase velocities are not the same. As a result, QNMs whose asymptotic phase velocity is $v_{p}\neq v_{g}$ in $(r,t)$ coordinates will undergo phase divergences at the boundaries. This can be removed by a phase-rotation of the field,
@@
where we introduce $\Delta=i(v_{g}-v_{p})/2$, so that the phase rotation is parameterized by both $v_{g}$ and $v_{p}$. The form of the required phase rotation follows from the asymptotic dispersion relation of Eq. (1) and the choice of height function, $h(x)$. Intuitively, it depends on the mismatch of the two velocities. The result is that the field $\hat{\phi}$ is bounded and well-defined on the new space.
The cost of the above construction is that we introduce two unknown real parameters, $v_{g}$ and $v_{p}$, into the problem. In fact, identifying velocity pairs that correspond to actual QNM solutions is as difficult a problem as determining the QNM spectrum itself. This may be seen by the relation,
$\Omega=-\frac{1}{2}v_{g}\bigg{(}v_{p}+i\sqrt{v_{g}(v_{g}-2v_{p})}\bigg{)},$ (5)
which we derive, in the Appendix, from the asymptotic dispersion relation of Eq. (1). This holds true for any mode whose asymptotic group and phase velocities are $v_{g}$ and $v_{p}$, respectively. The existence of a relation such as Eq. (5) is a direct consequence of dispersion. In a relativistic system, all asymptotic speeds are the speed of light, so $\Omega$ cannot be expressed in terms of asymptotic velocities. This difference between the relativistic and non-relativistic methods is crucial. Eqs. (3) and (4) mean we obtain an equation of motion, and an eigenvalue equation for the complex frequency $\Omega$, both of which are parameterized by $v_{g}$ and $v_{p}$. The significance of Eq. (5) is that these additional parameters can ultimately be eliminated, leaving $\Omega$ as the only unknown in the problem.
@@
where $J_{1}$ and $J_{2}$, given in the Appendix, are spatial operators depending on the potential and the asymptotic velocities. In contrast to the relativistic method, $\partial_{\tau}\hat{\psi}$ cannot be isolated by division in Eq. (6) because its pre-factor vanishes at $x=0$. This occurs because the contours of $\tau$ have to “turn around” in order to be outgoing in the $(r,t)$ coordinates. Our alternative approach is to construct $\partial_{\tau}\hat{\psi}$ using a Taylor series around $x=0$, obtaining the required derivatives by repeated differentiation of Eq. (6). In fact, this treatment is necessary only for terms indivisible by $x^{2}$, and we obtain a simpler result if we initially separate the terms in this way. This separation is mostly trivial, but for the potential, where we write $V=V_{0}+xV_{1}+x^{2}\tilde{V}(x)$, with $V_{0}$ and $V_{1}$ Taylor series coefficients about $x=0$ and $\tilde{V}$ accounting for the remaining terms. We obtain
@@
where $L_{1}$ and $L_{2}$ are spatial operators that we derive in the Appendix. Eq.(7) is formally identical to that obtained in the relativistic method Jaramillo2021 , but the operators are quite different, containing arbitrarily high spatial derivatives and depending on the asymptotic velocities.
@@
The operator $L$ is parameterized by $v_{g}$ and $v_{p}$, giving rise a family of operators. For each operator, Eq. (8) defines a unique eigenvalue problem and a corresponding spectrum. However, only a subset of the frequencies from these spectra obey Eq. (5), and it is this subset which comprises the QNM spectrum of Eq. (1). Using Eq. (5) to eliminate $v_{g}$ and $v_{p}$, one obtains a problem in which the frequency $\Omega$ is the only unknown and all solutions correspond to QNMs. In this formulation, $L$ is parameterized by $\Omega$, which constitutes an essential difference from the relativistic method, wherein the corresponding operator does not depend on $\Omega$ Jaramillo2021 . Importantly, Eq.(8) unambiguously determines the QNM spectrum.
@@
The QNM spectrum may then be obtained from Eq. (9) in the usual way using $\det(L^{N}-\Omega\ \!\text{Id})=0$. In the Appendix, we show that this determinant may be rewritten as that of a smaller $N$-dimensional matrix, $M$. Its elements are quadratic in the square root of the QNM frequency, giving rise to a polynomial of degree $2N$ in $\sqrt{\Omega}$. For a given potential $V$, the roots may be numerically determined in order to give $2N$ of the QNM frequencies. The fact that the frequency enters via its square root is a result of the Schrödinger equation having a first derivative in time, rather than a second derivative in time like the wave equation. Indeed, the exact QNM spectra of the Schrödinger and wave equations are related to each other by $i\sqrt{\Omega}=\omega$, as was elaborated in Burgess2024 . This means we can relate the results of the non-relativistic method to those of the relativistic method, allowing us to better evaluate the accuracy of the new method.
@@ III Quasinormal Modes of the Pöschl-Teller Potential
which serves as an exemplar for both the relativistic and non-relativistic methods. The QNMs of Eq. (10) are finite polynomials in the compactified spatial coordinate, with the result that an $N$-point discretization reproduces the first $2N$ QNMs to arbitrary precision. The Pöschl-Teller potential is also ideal because the corresponding QNM spectrum of the Schrödinger equation is given analytically by
@@
allowing us to verify our results Burgess2024 . In regards to the non-relativistic method, we note that the Pöschl-Teller potential is especially simple because all its QNMs have the same $v_{g}$ parameter, which is a result of the fact that $i\sqrt{\Omega}$ is aligned along vertical lines in the complex plane for this potential. While this simplicity does not influence the operation of the method, it does allow us to more easily assess the spectrum. Lastly, we partition the Pöschl-Teller potential with $V_{1}=0$ and $\tilde{V}=-V_{0}$, which reinforces the simplicity of the potential.
Figure 1: QNM spectra for the Schrödinger equation with a potential $V=V_{0}\hskip 1.5pt\text{sech}^{2}(r)+\epsilon\Delta V$, where $V_{0}=1$ and $\Delta V$ is the perturbation shown in green in the inset, alongside the unperturbed Pöschl-Teller potential in blue. The red dots and blue boxes correspond to the unperturbed Pöschl-Teller potential ($\epsilon=0$), with red dots () given by the exact formula and blue boxes ($\square$) numerically determined by the new method. The green crosses ($\times$) correspond to a perturbed potential ($\epsilon=10^{-30}$) and are also numerically determined. (a) displays the three QNM spectra, while (b) displays the same spectra under the transformation $\Omega\to i\sqrt{\Omega}$, which relates the spectra to those of a corresponding relativistic operator.
@@
In Figure 1(a), we plot the exact QNM frequencies of the unperturbed Pöschl-Teller potential, given in Eq. (11) Burgess2024 alongside those calculated by the new numerical method, with a resolution of $N=201$. We find excellent agreement for all frequencies, with an error which may be made arbitrarily small by increasing the working precision. These results are given in Figure 2. We also calculate the QNM spectrum of a perturbed Pöschl-Teller potential,
@@
where $\epsilon=10^{-30}$ and $\Delta V$ is a randomly chosen polynomial of degree 9, shown in the inset of Figure 1. We find that the spectrum for Eq.(12) closely resembles the unperturbed spectrum up to the $10^{\text{th}}$ overtone index, beyond which the frequencies are significantly displaced from their unperturbed values, as shown in Figure 2. These numerical results are then indicative of spectral instabilities that have been reported by previous authors Nollert1996 ; Daghigh2020 ; Jaramillo2021 .
Figure 2: Comparisons of exact and numerically determined QNM frequencies for the Schrödinger equation with a potential $V=V_{0}\hskip 1.5pt\text{sech}^{2}(r)+\epsilon\Delta V$, where $V_{0}=1$ and $\Delta V$ is the perturbation shown in Figure 1. The first 21 QNM frequencies are displayed. The unperturbed ($\epsilon=0$) spectrum is recovered well by the new numerical method, with errors smaller than $10^{-37}$ for the chosen working precision. We also obtain the perturbed ($\epsilon=10^{-30}$) spectrum and find the deviation from the exact spectrum grows rapidly with overtone index, $n$, as in previous works on spectral instability.
@@ IV Discussion
In this section, we discuss potential applications of the non-relativistic compactified hyperboloidal method that we developed in the preceding text, suggesting well-motivated directions in which to further develop the method and providing a sketch of how this can be achieved. The main motivations for this method were the modelling of QNMs of optical solitons, and the development of a framework within which one can treat QNMs in quantum gravity models with dispersive gravitational wave propagation. Beyond these, we note that this non-relativistic method may be employed equally well in any system governed by a Schrödinger equation equipped with a general potential. In this paper, we numerically calculated QNM spectra for the Pöschl-Teller potential and perturbations of that potential, finding agreement with earlier works Boonserm2011 ; Cardona2017 ; Burgess2024 . For potentials with different long-range behaviour than the Pöschl-Teller potential, one typically requires different choices of height function $h(x)$, but this requirement is shared by the relativistic method, and may be addressed by the same techniques Zenginoglu2011 ; Jasiulek2011 ; Macedo2018 . In addition, we note that this method may also be used to numerically solve for the quantum mechanical bound-states of a general potential well, using the well-known connection between the QNMs of a potential barrier and the bound-states of the corresponding well Ferrari1984a ; Ferrari1984b ; Churilova2022 ; Volkel2022 .
As described above, the non-relativistic method we have presented is closely related to the relativistic method, sharing many essential features with it. For instance, the classes of potentials that can be treated by the two methods are the same, and they have the same maximum achievable accuracy for a given resolution. As a result, the methods are comparable in their scope and power. They also share the same advantages and disadvantages when compared to other popular numerical methods, such as Leaver’s continued fraction method Leaver1986 . For example, in this case, both the relativistic and non-relativistic methods enjoy the advantage that they recover the entire spectrum simultaneously, and do not require initial seed values close to the QNM frequencies one wishes to compute Leaver1986 ; Zenginoglu2009 ; Ansorg2016 ; Jaramillo2021 .
@@
A further motivation for generalizing the non-relativistic method is to shed light on QNM spectral instabilities, and facilitate experimental tests of the recent ultraviolet universality conjecture, which posits that sufficiently high overtones converge to logarithmic Regge branches in the complex plane, in the high-frequency limit of potential perturbations Jaramillo2021 ; Boyanov2024 . This effect is easily seen in numerical calculations of the Pöschl-Teller spectrum, on account of its simplicity, but has yet to be experimentally confirmed. Using the optical soliton, whose perturbations realize this potential, experimental tests become possible. The numerical method presented above is essential for the modelling of these experiments, as one cannot realize an exact soliton in practice, and must always work with near-soliton potentials. In addition, higher-order dispersive effects will also be present in any experiment, and these must be understood in order to interpret observations of QNM spectral migration with the soliton. In particular, the influence of weak third-order dispersion acting on the perturbative probe field should be incorporated into the analysis, in order to provide the best test of the above conjecture. This motivates the development of the non-relativistic method beyond the Schrödinger equation, to include higher-order dispersive terms.
@@
with $\alpha(z)$ and $\beta(z)$ finite polynomials in $z$, and $d$ the larger degree among the two polynomials. In principle, we can apply a hyperboloidal coordinate transformation and a phase rotation of the fields, parameterised by the asymptotic velocities, $v_{g}$ and $v_{p}$. Then, we introduce auxiliary fields to effect a $d$th-order reduction in time, defining
@@
with $1\leq k<d$. Eq. (14) closely mirrors the treatment of resonator QNMs in optics lalanne2019 . The equations of motion of these fields are trivial for all fields but $\phi_{d}$, whose equation of motion more closely resembles Eq. (6). If we use a Taylor series expansion of $\partial_{\tau}\phi_{d}$ around zero, we can write it in terms of spatial operators acting on the fields. The general form of the now $d$-dimensional operator $L$ is
@@
which we discretize as before. Then, we use the asymptotic dispersion relation of Eq. (13) to eliminate the asymptotic velocities, obtaining a vector equation for the QNM frequencies. From Eq. (LABEL:LMatrix), it can be shown that it is always possible to construct an $N$-dimensional matrix $M$ whose determinant is a finite polynomial for the QNM frequencies. This may then be solved numerically and the frequencies $\Omega$ determined. This generalization is largely straight-forward. However, the divergences in space are multi-exponential with higher derivatives, leading to non-polynomial modes in the compactified coordinates. This complicates the imposition of QNM boundary conditions, and further work is required to address this. For example, approaches that augment the function space to include additional non-polynomial functions can be investigated. Future work can investigate how this generalized method compares with other numerical schemes, as the connection to the relativistic method is less concrete in this case.
The method presented is primarily intended for the gravitational context and long-range potentials, but the authors note that extensions to optical cavities or plasmonic resonators may be possible. Beyond QNMs, the non-relativistic method can be applied to spectra of non-selfadjoint operators, connecting with a larger research effort. We believe an explicit formulation in this context is a promising research direction. In addition, future works can develop the method, along the lines of Jaramillo2021 , in order to calculate the pseudospectra of non-relativistic operators. It is our view that the relationship between perturbed QNM spectra and the pseudospectrum is best understood from a broader perspective, not limited to relativistic wave operators. We expect that numerical methods will become increasingly important for addressing questions in the theory of QNMs, and anticipate that investigations into the QNMs of non-relativistic fields will provide new avenues to explore these questions.
@@
This work was supported in part by the Science and Technology Facilities Council through the UKRI Quantum Technologies for Fundamental Physics Programme [Grant ST/T005866/1]. CB was supported by the UK Engineering and Physical Sciences Research Council [Grant EP/T518062/1].
We would like to express our thanks to Théo Torres for providing us a useful overview at the outset of this research.